OpenStudy (anonymous):

Can someone plz tell me how to do 48 divided by 0.012 withoutusing a calculator ta

OpenStudy (solomonzelman):

well, \(\large\color{black}{ 48\div 0.012= }\) \(\large\color{black}{ 48\div 12\div\frac{\Large1}{\Large1000} }\)

OpenStudy (solomonzelman):

first divide by 48 by 12, (you know what that will be, and then apply the rule \(\large\color{black}{ A\div\frac{\Large V}{\Large Z}~~=>~~A\div\frac{\Large Z}{\Large V} }\)

OpenStudy (solomonzelman):

If there is something you are uncertain about please ask...

OpenStudy (anonymous):

Still dont gt it but thanks

OpenStudy (solomonzelman):

well, can you do the first step? We know that 48 divided by 12 is 4.

OpenStudy (solomonzelman):

well, can you do the first step? We know that 48 divided by 12 is 4.

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

My latex isn't working for me.

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ 4 \div {\LARGE \frac{1}{1000}} }\) \(\LARGE\color{blue}{ 4 \times {\LARGE \frac{1000}{1}} }\) \(\LARGE\color{blue}{ 4000 }\)

OpenStudy (anonymous):

Thanks...

OpenStudy (anonymous):

Could you plz tell me how to do 15 x 0.04 without a calc

OpenStudy (solomonzelman):

15 divided by 4 divided by 1/100

OpenStudy (solomonzelman):

15 divided by 4 divided by 1/100 15 divided by 1/100 divided by 4 15 times 100 divided by 4 1500 divided by 4 =375 4 1500 12 -- 30 28 -- 20 20 -- 0

OpenStudy (solomonzelman):

375

OpenStudy (anonymous):